# U-REPA: Aligning Diffusion U-Nets to ViTs

### v2

@@ 1 Introduction
Representation Alignment (REPA) [repa], a methodology that aligns features from Diffusion Transformers (DiT) [dit] to modern visual encoders, has been demonstrated to significantly accelerate DiT training. This approach holds particular significance given the growing prominence of DiTs, which have gained mainstream adoption in diffusion models and are extensively applied across image generation [pixartalpha; sd3; flux] and video generation domains [opensora; opensoraplan; pyramidflow]. However, emerging empirical evidence suggests that U-Net [unet] architectures might present a more advantageous alternative to DiTs in certain scenarios [simplediffusion; hourglass; udit; dic]: U-Net-based models exhibit substantially faster convergence while achieving generation quality comparable to their transformer-based counterparts. This dichotomy motivates our core research inquiry - can modern Vision Transformer (ViT [vit])-based visual encoders be effectively adapted to guide diffusion U-Net training through alignment mechanisms similar to REPA, thereby potentially elevating the convergence speed ceiling of diffusion models?
@@ 2 Related Work
The development of diffusion architectures. The conventional diffusion works [ddpm; ddim; songunet; dhariwal] leverages a U-Net [unet] architecture, whose basic block is a concatenation of convolution layers and self-attention. More recent architectural innovations in diffusion models have witnessed a paradigm shift from conventional U-Net frameworks toward transformer-based architectures. The emergence of Diffusion Transformers [uvit; dit] demonstrates their competitive performance despite abandoning the inductive biases inherent in U-Net designs. U-ViT [uvit] represents an intermediate architecture that preserves U-Net’s hierarchical structure but replaces convolutional blocks with transformer layers, notably omitting the traditional downsampling operations. Subsequent developments have further streamlined the architecture: DiT [dit] adopts a pure transformer backbone with isotropic scaling, while SiT [sit] integrates the transformer architecture into the RectifiedFlow framework. Some other works either improve the micro-designs [visionllama; fit], or focuses on architectural efficiency [pixartsigma; sana; lit].
In contrary to these DiT works, some works still sticks to U-Net architectures and offer valuable rethinking on this conventional architectural preference: in pixel-space image generation, works including SimpleDiffusion [simplediffusion] and HourglassDiT [hourglass] still sticks to U-Net; in with its variants like U-DiT [udit], Playground v3 [playgroundv3], and DiC [dic] extending its success to latent-space diffusion through simple Conv33 designs. While these implementations empirically validate U-Net’s accelerated convergence and stable training dynamics compared to transformer-based alternatives, current research predominantly focuses on proposing architectural modifications rather than uncovering the reasons of U-Net’s superior diffusion performance.
Techniques for better DiT performance. Building upon the success of self-supervised learning [mae], MDT [mdt] and MaskDiT [maskdit] pioneer masked image modeling in diffusion frameworks by adaptively masking a good proportion of input patches during training. Other than the masking strategy, a bunch of diffusion works refer to higher-level semantic guidance from off-the-shelf pretrained models that significantly improves generation quality. REPA [repa] establishes feature alignment between ViT-based encoder embeddings and diffusion latent spaces through contrastive learning. LightningDiT [vavae] innovates through an improved VAE distilled from MAE [mae] and DINOv2 [dinov2]. Ma et al. [diffcot] introduces CLIP [clip] and DINO [dino] to verify inference-time scaling of diffusion models. These methods demonstrate that a higher-level semantic-rich feature-map from pretrained vision encoders is helpful to diffusion-based generation.
@@ 3.1 Preliminaries: REPA for DiT
Representation Alignment (REPA) [repa] distills Diffusion Transformers with semantic features from off-the-shelf ViT-based vision encoders (e.g. DINOv2 [dinov2], CLIP [clip], MAE [mae], et cetera). Given a ViT-based vision encoder and clean image , let denote its patch embeddings, where and represent the number of patches and embedding dimension, respectively. REPA establishes feature alignment between and the projected diffusion encoder outputs , where is the latent representation from the diffusion transformer at timestep , and is a trainable multilayer perceptron (MLP).
@@
where denotes a similarity metric (e.g., cosine similarity). Typically, is adopted as the output from early layers (the original work adopts layer index 8) in DiT for better alignment. This alignment term is combined with the basic flow-based diffusion objective (i.e. SiT [sit]) through a tunable coefficient , and the final loss for diffusion model training is formulated as follows:
@@ 3.2 Evaluating the Potential of U-Net
In diffusion models, U-Net and isotropic architectures (e.g., DiT) exhibit distinct design philosophies. While DiT achieves state-of-the-art results through scalability and integration with advanced techniques, U-Net-based methods emphasize faster convergence [udit]. To dissect U-Net’s efficacy, we isolate its two core components: skip connections and downsampling.
@@
Building on this insight, we propose DiT (and SiT for the flow-based version, following the naming convention of [sit]) by adding tricks of RoPE [rope] and SwiGLU following previous work [visionllama; udit]
@@ 3.3 Aligning U-Net to ViT Encoders
Position for alignment. Regarding the concern that block functionalities differ between U-Net and DiT, the comparison of prior studies [diffpruning; deltadit] demonstrates divergent hierarchical specialization: U-Net architectures typically employ mid-network layers for high-level semantic synthesis while reserving shallow layers for low-level image refinement, whereas DiTs exhibit a totally different pattern - early layers primarily govern semantic-rich outline formation with deeper layers handling detailed image refinement.
@@
Some recent works on Diffusion U-Net [unified; freeu] reveals that higher-stage U-Net features are low-frequency subspaces that discards higher frequency compoenents, including noises. Gaps are inevitable when evaluating the cosine similarities of detail-rich, high-frequency-rich vectors and flat, low-frequency-dominated vectors. In this sense, strict token-level alignment constraints like the original REPA loss prove suboptimal under these conditions, as they assume implicit feature space homogeneity between aligned modalities.
@@
Rather than strict token-wise regularization, we resort to looser objectives that does not require rigid augular alignment. Inspired by manifold knowledge distillation [manifoldKD], we hold that aligning similarities between samples from the same feature space could be a promising solution. Hence, we define Manifold Loss as
@@ 4.1 Experiment Setup
Experiment settings. Our implementation completely adheres to the training protocol established in REPA [repa]. Following the architectural configuration of latent diffusion models [sd], we employ the identical VAE variant (sd-vae-ft-ema) and adopt the AdamW optimizer. To ensure fair comparison, we maintain identical hyperparameter settings across all experiments: a global batch size of 256, fixed learning rate of , and disabled weight decay (set to 0). () is set as (0.9, 0.999). All experiments are conducted on the ImageNet 2012 benchmark [imagenet] under a controlled environment with a fixed random seed (global seed=0). 8 NVIDIA A100 GPUs are used for main experiments.
For main experiments (Tab. 8), we apply guidance interval [guidanceinterval] and SDE sampling according to the convention of REPA [repa] for fair comparison. We select smaller cfg of 1.65, because we found it is better for our architecture, different from SiT. For all ablation experiments, we train models for 100K iterations, which is sufficient to show the trend of model performance; sampling is conducted with the default setting of the official REPA codebase, i.e. in ODE and guidance interval .
@@ 4.3 Ablation Studies
| ADM-U [dhariwal] | 400 | 3.94 |
| VDM [vdmpp] | 560 | 2.40 |
| Simple diffusion [simplediffusion] | 800 | 2.77 |
@@ Appendix A Other Improvements
Time-aware MLP. Some works on in the diffusion task [dtr, dynamicdit] reveals that a channel dimension is particularly sensitive and useful to a certain subset of time during sampling. As the time-variant feature is to be aligned to time-invariant vision encoder features, we hold that alignment could perform better when the MLP is time-aware and extract time-invariant information out of diffusion features for alignment.
Inspired by the conditioning of canonical Diffusion U-Net [ddpm, songunet, dhariwal] and DiT [dit], we add a module to predict a pair of channel-wise shift& scale vector . The module is in parallel to MLP and follows the design of DiT’s AdaLN, which is a concatenation of a SiLU and a Linear layer. The shift& scale vectors are imposed on the output MLP as follows:
Weight schedule of loss. As is prompted in REPA [repa], designing weight schedule is a future direction. We try various weight schedules (i.e. make in Eq. 2 a function with respect to time) but found that these schedules bring very limited improvements on the proposed U-REPA. Hence, we stick to the original constant weight strategy of REPA.
@@ Appendix C Limitations & Impact
Limitations and Future work. The U-Net architecture is a simple one with only one intermediate stage. We do not further refine the architecture as we want to show U-Net architectures as simple as SiT could also achieve rapid convergence. Further improvements on the U-Net architecture includes efficient attention [sana, udit], non-integer down& up scaling factors [fdvit], and more use of convolutions [songunet, dic]. Besides, whether U-REPA could be applied to downstream diffusion tasks that rely heavily on U-Nets (e.g. Low-Level Vision) remains to be investigated.

### v3

@@ 1 Introduction
Representation Alignment (REPA) [45], a methodology that aligns features from Diffusion Transformers (DiT) [30] to modern visual encoders, has been demonstrated to significantly accelerate DiT training. This approach holds particular significance given the growing prominence of DiTs, which have gained mainstream adoption in diffusion models and are extensively applied across image generation [4; 11; 23] and video generation domains [48; 22; 19]. However, emerging empirical evidence suggests that U-Net [33] architectures might present a more advantageous alternative to DiTs in certain scenarios [18; 7; 39; 38]: U-Net-based models exhibit substantially faster convergence while achieving generation quality comparable to their transformer-based counterparts. This dichotomy motivates our core research inquiry - can modern Vision Transformer (ViT [10])-based visual encoders be effectively adapted to guide diffusion U-Net training through alignment mechanisms similar to REPA, thereby potentially elevating the convergence speed ceiling of diffusion models?
@@ 2 Related Work
The development of diffusion architectures. The conventional diffusion works [17; 35; 36; 9] leverages a U-Net [33] architecture, whose basic block is a concatenation of convolution layers and self-attention. More recent architectural innovations in diffusion models have witnessed a paradigm shift from conventional U-Net frameworks toward transformer-based architectures. The emergence of Diffusion Transformers [1; 30] demonstrates their competitive performance despite abandoning the inductive biases inherent in U-Net designs. U-ViT [1] represents an intermediate architecture that preserves U-Net’s hierarchical structure but replaces convolutional blocks with transformer layers, notably omitting the traditional downsampling operations. Subsequent developments have further streamlined the architecture: DiT [30] adopts a pure transformer backbone with isotropic scaling, while SiT [26] integrates the transformer architecture into the RectifiedFlow framework. Some other works either improve the micro-designs [6; 25], or focuses on architectural efficiency [3; 42; 40].
In contrary to these DiT works, some works still sticks to U-Net architectures and offer valuable rethinking on this conventional architectural preference: in pixel-space image generation, works including SimpleDiffusion [18] and HourglassDiT [7] still sticks to U-Net; in with its variants like U-DiT [39], Playground v3 [24], and DiC [38] extending its success to latent-space diffusion through simple Conv33 designs. While these implementations empirically validate U-Net’s accelerated convergence and stable training dynamics compared to transformer-based alternatives, current research predominantly focuses on proposing architectural modifications rather than uncovering the reasons of U-Net’s superior diffusion performance.
Techniques for better DiT performance. Building upon the success of self-supervised learning [16], MDT [13] and MaskDiT [47] pioneer masked image modeling in diffusion frameworks by adaptively masking a good proportion of input patches during training. Other than the masking strategy, a bunch of diffusion works refer to higher-level semantic guidance from off-the-shelf pretrained models that significantly improves generation quality. REPA [45] establishes feature alignment between ViT-based encoder embeddings and diffusion latent spaces through contrastive learning. LightningDiT [44] innovates through an improved VAE distilled from MAE [16] and DINOv2 [28]. Ma et al. [27] introduces CLIP [31] and DINO [2] to verify inference-time scaling of diffusion models. These methods demonstrate that a higher-level semantic-rich feature-map from pretrained vision encoders is helpful to diffusion-based generation.
@@ 3.1 Preliminaries: REPA for DiT
Representation Alignment (REPA) [45] distills Diffusion Transformers with semantic features from off-the-shelf ViT-based vision encoders (e.g. DINOv2 [28], CLIP [31], MAE [16], et cetera). Given a ViT-based vision encoder and clean image , let denote its patch embeddings, where and represent the number of patches and embedding dimension, respectively. REPA establishes feature alignment between and the projected diffusion encoder outputs , where is the latent representation from the diffusion transformer at timestep , and is a trainable multilayer perceptron (MLP).
@@
where denotes a similarity metric (e.g., cosine similarity). Typically, is adopted as the output from early layers (the original work adopts layer index 8) in DiT for better alignment. This alignment term is combined with the basic flow-based diffusion objective (i.e. SiT [26]) through a tunable coefficient , and the final loss for diffusion model training is formulated as follows:
@@ 3.2 Evaluating the Potential of U-Net
In diffusion models, U-Net and isotropic architectures (e.g., DiT) exhibit distinct design philosophies. While DiT achieves state-of-the-art results through scalability and integration with advanced techniques, U-Net-based methods emphasize faster convergence [39]. To dissect U-Net’s efficacy, we isolate its two core components: skip connections and downsampling.
@@ 3.3 Aligning U-Net to ViT Encoders
Position for alignment. Regarding the concern that block functionalities differ between U-Net and DiT, the comparison of prior studies [12; 5] demonstrates divergent hierarchical specialization: U-Net architectures typically employ mid-network layers for high-level semantic synthesis while reserving shallow layers for low-level image refinement, whereas DiTs exhibit a totally different pattern - early layers primarily govern semantic-rich outline formation with deeper layers handling detailed image refinement.
@@
Some recent works on Diffusion U-Net [41; 34] reveals that higher-stage U-Net features are low-frequency subspaces that discards higher frequency compoenents, including noises. Gaps are inevitable when evaluating the cosine similarities of detail-rich, high-frequency-rich vectors and flat, low-frequency-dominated vectors. In this sense, strict token-level alignment constraints like the original REPA loss prove suboptimal under these conditions, as they assume implicit feature space homogeneity between aligned modalities.
@@
Rather than strict token-wise regularization, we resort to looser objectives that does not require rigid augular alignment. Inspired by manifold knowledge distillation [14], we hold that aligning similarities between samples from the same feature space could be a promising solution. Hence, we define Manifold Loss as
@@ 4.1 Experiment Setup
Experiment settings. Our implementation completely adheres to the training protocol established in REPA [45]. Following the architectural configuration of latent diffusion models [32], we employ the identical VAE variant (sd-vae-ft-ema) and adopt the AdamW optimizer. To ensure fair comparison, we maintain identical hyperparameter settings across all experiments: a global batch size of 256, fixed learning rate of , and disabled weight decay (set to 0). () is set as (0.9, 0.999). All experiments are conducted on the ImageNet 2012 benchmark [8] under a controlled environment with a fixed random seed (global seed=0). 8 NVIDIA A100 GPUs are used for main experiments.
For main experiments (Tab. 8), we apply guidance interval [21] and SDE sampling according to the convention of REPA [45] for fair comparison. We select smaller cfg of 1.65, because we found it is better for our architecture, different from SiT. For all ablation experiments, we train models for 100K iterations, which is sufficient to show the trend of model performance; sampling is conducted with the default setting of the official REPA codebase, i.e. in ODE and guidance interval .
@@ 4.3 Ablation Studies
| ADM-U [9] | 400 | 3.94 |
| VDM [20] | 560 | 2.40 |
| Simple diffusion [18] | 800 | 2.77 |
@@ Appendix A Other Improvements
Time-aware MLP. Some works on in the diffusion task [29, 46] reveals that a channel dimension is particularly sensitive and useful to a certain subset of time during sampling. As the time-variant feature is to be aligned to time-invariant vision encoder features, we hold that alignment could perform better when the MLP is time-aware and extract time-invariant information out of diffusion features for alignment.
Inspired by the conditioning of canonical Diffusion U-Net [17, 36, 9] and DiT [30], we add a module to predict a pair of channel-wise shift& scale vector . The module is in parallel to MLP and follows the design of DiT’s AdaLN, which is a concatenation of a SiLU and a Linear layer. The shift& scale vectors are imposed on the output MLP as follows:
Weight schedule of loss. As is prompted in REPA [45], designing weight schedule is a future direction. We try various weight schedules (i.e. make in Eq. 2 a function with respect to time) but found that these schedules bring very limited improvements on the proposed U-REPA. Hence, we stick to the original constant weight strategy of REPA.
@@ Appendix C Limitations & Impact
Limitations and Future work. The U-Net architecture is a simple one with only one intermediate stage. We do not further refine the architecture as we want to show U-Net architectures as simple as SiT could also achieve rapid convergence. Further improvements on the U-Net architecture includes efficient attention [42, 39], non-integer down& up scaling factors [43], and more use of convolutions [36, 38]. Besides, whether U-REPA could be applied to downstream diffusion tasks that rely heavily on U-Nets (e.g. Low-Level Vision) remains to be investigated.